OpenStudy (anonymous):

On his university application, Prashad must list his course choices in order of preference. He must choose four of the six courses available in his major discipline and three of the four courses offered in related subjects. In how many ways can Prashad list his course choices? Explain the reasoning for your answer.

OpenStudy (anonymous):

Since the wording of this question says "how many ways can he list his courses" the order in which they are picked does matter right?

OpenStudy (anonymous):

So I use the permutation 6*5*4*3=360 For his major disciplines

OpenStudy (anonymous):

three of the four courses=4*3*2=24 Therefore the order in which he can list the course equals to 360*24=8640

OpenStudy (anonymous):

@alekos Please

OpenStudy (anonymous):

Can someone confirm that my solutions are right?

OpenStudy (anonymous):

@campbell_st

OpenStudy (anonymous):

@UsukiDoll

OpenStudy (anonymous):

What do you think?

OpenStudy (campbell_st):

here is my best guess  multiply the choices for major and minor together for the answer